A bookself has space for 6 books. If 8 different books are available, how many different arrangements can be made on the shelf?
My answer was 28. I found this by using the following combination:
(8 nCr 6) = 28
I was marked wrong, what is the correct way to do this problem and how can i avoid this mistake in the future?

Answer by stanbon(75887)

A bookself has space for 6 books. If 8 different books are available, how many different arrangements can be made on the shelf?
---
Ans: 8P6 = 8!/(8-6)! = 8*7*6*5*4*3 = 20160
============================================
Cheers,
Stan H.
===============